OpenStudy (anonymous):

a guy wire from the top of a 20-foot pole is anchored to the ground 12 ft from the base of the pole. how long is the wire?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

you can use the pythagorean theorem to find z a^2 + b^2 = c^2 20^2 + 12^2 = z^2 z = ??

OpenStudy (anonymous):

544=z^2

OpenStudy (anonymous):

but thats not a perfect square

jimthompson5910 (jim_thompson5910):

it usually isn't but it still works because you can have fractional lengths (like 2.3 or something)

jimthompson5910 (jim_thompson5910):

Take the square root of both sides. What do you get?

OpenStudy (anonymous):

23.323..

jimthompson5910 (jim_thompson5910):

So that is the approximate length of z

jimthompson5910 (jim_thompson5910):

which is the approximate length of the wire in feet

OpenStudy (anonymous):

oh ok thanks

jimthompson5910 (jim_thompson5910):

you're welcome